OpenStudy (anonymous):

x plus 1/5 of y is 14, and sum of 1/2 of x and i/5 of y is 5 Solve for x and y

OpenStudy (anonymous):

1 x + --y = 14 ---------(1) 5 1 1 --x + --y = 5 ---------(2) 2 5 multiply (1) throughout by 5 and (2) by 10 5x + y = 70 --------(3) 5x + 2y = 50 --------(4) subtract (4) from (3), to get -y = 20 y = -20 now substitute y = -20 in (3) 5x + (-20) = 70 5x - 20 = 70 5x = 70 + 20 x = 90/5 x = 18 so solution (x,y) = (18, -20)

OpenStudy (anonymous):

Hermph thats why I got the answer wrong...Thanks.

OpenStudy (anonymous):

whenever u hv fractions, follow the above procedure to get rid of them....☺

OpenStudy (anonymous):

:)